OpenStudy (anonymous):

the length of a rectangle is three times the width. express the area, A as a function of the length, L

OpenStudy (anonymous):

let the width be x. then, the length will be 3x=L. therefore width x = L/3 Area = length * width A= L * x A = Lx or A = L*L/3 (since x= L/3) A = L^2 /3

OpenStudy (saifoo.khan):

@shubham.bagrecha is correct. (:

OpenStudy (anonymous):

A=L*W Since the lengh is 3 times the width then Area=L*1/3L